OpenStudy (anonymous):

Use f(x) =x+4 and g(x)=x+2 to find each of the following? A. (f+g) (x) B. (f-g) (x) C. (fg) (x) D. (f/g) (x) A. (f+x) (x)=???

OpenStudy (anonymous):

I'm sorry I meant A. (f+g) (x) ??

OpenStudy (harsimran_hs4):

(f+g)x = f(x) + g(x) sum of the individual values of the functions

OpenStudy (anonymous):

\[ (f+g) (x) = f(x) + g(x) = (x+4) +(x+2) \]

OpenStudy (harsimran_hs4):

@Virgo91684 clear??

OpenStudy (anonymous):

Yes and how do I simplify my answers?

OpenStudy (anonymous):

I was able to get all four but my homework is asking for a simplified answer!

OpenStudy (harsimran_hs4):

(f+g) x = 2x + 6 you cannot simplify this expression more... it`s already in the simplified form